OpenStudy (anonymous):

Amy and Elizabeth both work in sales. Each week Amy earns $18/h plus a $40 bonus, and Elizabeth earns $12/h plus a $100 bonus. How many hours must Amy and Elizabeth each work in a week in order to earn the same amount of money? For that many hours, how much money would each earn? They would each have to work__ hours, and they would earn $ ___.

OpenStudy (anonymous):

plz help i tried 5 times and been doing this for 3 hr plz plz help

OpenStudy (anonymous):

Hey there :) You could try trial and error but I did it using two sides of an equation: So if they both have to earn the same amount with the same amount of hours they could be put on opposite sides of an equation 18h + 40 = 12h + 100 Then you move all the terms with h in it to one side (18h -12h) + 40 = (12h -12h) + 100 6h + 40 = 100 Then move all the integers to the other 6h + (40 - 40) = (100 - 40) 6h = 60 Then divide both sides by 6 to have the h stand alone 6h/6 = 60/6 h = 10 So they both have to work 10 hours to make the same amount of money. Which will be $220, because (18 x 10) + 40 = 220 Hope this helps :)